+ d - 7 17. [1pt] A wheel of mass m = 1.87 kg and moment of inertia, 1= 1.530 kg.m is supported by a pivot a distance d=0.17Could you please help with question 18-20
Answer #1

18. Initial angular momentum = I \omega = 1.530 \times 90 \approx 138 kg m2 s-1

19. Torque = mg \times d = 1.87 \times 9.8 \times 0.17 = 3.115 N m

20. magnitude of angular velocity \omega p with which the wheel and axle precess

\omegap = mgd / ( I \omega ) = ( 1.87 \times 9.8 \times 0.17 ) / 138 = 0.0226 rad/s